# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

### v7

@@ Abstract
Training logistic regression over encrypted data has been a compelling approach in addressing security concerns for several years. In this paper, we introduce a more efficient gradient variant, called quadratic gradient, for privacy-preserving logistic regression training. We enhance Nesterov’s Accelerated Gradient (NAG) and Adaptive Gradient Algorithm (Adagrad) by incorporating their quadratic gradients and evaluate these improved algorithms on various datasets. Experiments show that the enhanced algorithms have state-of-the-art performance in terms of convergence speed when compared to their raw first-order gradient methods. Also, we employ the enhanced NAG method to implement homomorphic logistic regression training, achieving comparable results in only 4 iterations. There is a promising possibility that quadratic gradient could combine first-order gradient descent/ascent algorithms with the second-order Newton-Raphson method, making it applicable to a wide range of numerical optimization problems.
Homomorphic Encryption, Logistic Regression, Quadratic Gradient, Simplified Fixed Hessian, Nesterov’s Accelerated Gradient, Gradient Ascent Algorithm
@@ 1 Introduction
Given a person’s healthcare data related to a certain disease, we can train a logistic regression (LR) model capable of telling whether or not this person is likely to develop this disease. However, such personal health information is highly private to individuals. The privacy concern, therefore, becomes a major obstacle for individuals to share their biomedical data. The most secure solution is to encrypt the data into ciphertexts first by Homomorphic Encryption (HE) and then securely outsource the ciphertexts to the cloud, without allowing the cloud to access the data directly. iDASH is an annual competition that aims to call for implementing interesting cryptographic schemes in a biological context. Since 2014, iDASH has included the theme of genomics and biomedical privacy. The third track of the 2017 iDASH competition and the second track of the 2018 iDASH competition were both to develop homomorphic-encryption-based solutions for building an LR model over encrypted data.
Several studies on logistic regression models are based on homomorphic encryption. Kim et al. (2018b) discussed the problem of performing LR training in an encrypted environment. They used the full batch gradient descent in the training process and the least-squares method to get the approximation of the sigmoid function. In the iDASH 2017 competition, Bonte and Vercauteren (2018), Kim et al. (2018a), Chen et al. (2018), and Crawford et al. (2018) all investigated the same problem that Kim et al. (2018b) studied. In the iDASH competition of 2018, Kim et al. (2019b) and Blatt et al. (2019) further worked on it for an efficient packing and semi-parallel algorithm. There are other related works (Kim et al., 2019a; Bergamaschi et al., 2019) focusing on various aspects but the papers most relevant to this work are (Bonte & Vercauteren, 2018) and (Kim et al., 2018a). Bonte and Vercauteren (2018) developed a practical algorithm called the simplified fixed Hessian (SFH) method. Our study extends their work and adopts the ciphertext packing technique proposed by Kim et al. (2018a) for efficient homomorphic computation.
@@
    We propose a new gradient variant, quadratic gradient, which can combine the first-order gradient algorithms and the second-order Newton-Raphson method (aka Newton’s method) as one.
@@
    We develop two enhanced gradient algorithms by equipping the original ones with quadratic gradient. The resulting algorithms exhibit state-of-the-art performance in terms of convergence speed.
@@
    We implement privacy-preserving logistic regression training using the enhanced NAG method, to our best knowledge, which seems to be a good choice without compromising much on computation and storage.
@@ 2 Preliminaries
We use the square brackets “[][\ ][ ]” to denote the index of a vector or matrix element in what follows. For example, for a vector $\boldsymbol{v}\in\mathbb{R}^{(n)}$ and a matrix $M\in\mathbb{R}^{m\times n}$, $\boldsymbol{v}[i]$ or $\boldsymbol{v}_{[i]}$ means the $i$-th element of vector $\boldsymbol{v}$ and $M[i][j]$ or $M_{[i][j]}$ the $j$-th element in the $i$-th row of $M$.
@@ 2.1 Fully Homomorphic Encryption
Fully Homomorphic Encryption (FHE) is a type of cryptographic scheme that can be used to compute an arbitrary number of additions and multiplications directly on the encrypted data. It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation (2009). FHE schemes themselves are computationally time-consuming; the choice of dataset encoding matters likewise to the efficiency. In addition to these two limits, how to manage the magnitude of plaintext (Jäschke & Armknecht, 2016) also contributes to the slowdown. Cheon et al. (2017) proposed a method to construct an HE scheme with a rescaling procedure which could eliminate this technical bottleneck effectively. We adopt their open-source implementation HEAAN while implementing our homomorphic LR algorithms. In addition, it is inevitable to pack a vector of multiple plaintexts into a single ciphertext for yielding a better amortized time of homomorphic computation. HEAAN supports a parallel technique (aka SIMD) to pack multiple complex numbers in a single polynomial and provides rotation operation on plaintext slots. The underlying HE scheme in HEAAN is well described in (Kim et al., 2018a, b; Han et al., 2019).
@@ 2.2 Database Encoding Method
Kim et al. (2018a) proposed an efficient and promising database-encoding method by using SIMD technique, which could make full use of the computation and storage resources. Suppose that a database has a training dataset consisting of $n$ samples with $(1+d)$ covariates, they packed the training dataset $Z$ into a single ciphertext in a row-by-row manner.
@@
Han et al. (2019) mentioned several basic but important operations used by Kim et al. (2018a) in their implementation, such as a procedure named “SumColVec” to compute the summation of the columns of a matrix. With these fundamental operations, more intricate computations, like computing gradients in logistic regression models, become achievable.
@@ 2.3 Logistic Regression Model
Logistic regression is widely used in binary classification tasks to infer whether a binary-valued variable belongs to a certain class or not. LR can be generalized from linear regression (Murphy, 2012) by mapping the whole real line $(\boldsymbol{\beta}^{\top}\mathbf{x})$ to $(0,1)$ via the sigmoid function $\sigma(z)=1/(1+\exp(-z))$, where the vector $\boldsymbol{\beta}\in\mathbb{R}^{(1+d)}$ is the main parameter of LR and the vector $\mathbf{x}=(1,x_{1},\ldots,x_{d})\in\mathbb{R}^{(1+d)}$ the input covariate. Thus logistic regression can be formulated with the class label $y\in\{\pm 1\}$ as follows:
$\Pr(y=+1|\mathbf{x},\boldsymbol{\beta})$ $=\sigma(\boldsymbol{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{-\boldsymbol{\beta}^{\top}\mathbf{x}}},$
$\Pr(y=-1|\mathbf{x},\boldsymbol{\beta})$ $=1-\sigma(\boldsymbol{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{+\boldsymbol{\beta}^{\top}\mathbf{x}}}.$
@@
The logistic regression problem can be transformed into an optimization problem that seeks a parameter $\boldsymbol{\beta}$ to maximize $L(\boldsymbol{\beta})=\prod_{i=1}^{n}\Pr(y_{i}|\mathbf{x}_{i},\boldsymbol{\beta})$ or its log-likelihood function $l(\boldsymbol{\beta})$ for convenience in the calculation:
$l(\boldsymbol{\beta})=\ln L(\boldsymbol{\beta})=-\sum_{i=1}^{n}\ln(1+e^{-y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i}}),$
where $n$ is the number of examples in the training dataset. LR does not have a closed form of maximizing $l(\boldsymbol{\beta})$ and two main methods are adopted to estimate the parameters of an LR model: (a) gradient descent method via the gradient; and (b) Newton’s method by the Hessian matrix. The gradient and Hessian of the log-likelihood function $l(\boldsymbol{\beta})$ are given by, respectively:
$\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})$ $=\sum_{i}(1-\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i}))y_{i}\mathbf{x}_{i},$
$\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$ $=\sum_{i}(y_{i}\mathbf{x}_{i})(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})(y_{i}\mathbf{x}_{i})$
@@
where $S$ is a diagonal matrix with entries $S_{ii}=(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})$ and $X$ the dataset.
The log-likelihood function $l(\boldsymbol{\beta})$ of LR has at most a unique global maximum (Allison, 2008), where its gradient is zero. Newton’s method is a second-order technique to numerically find the roots of a real-valued differentiable function, and thus can be used to solve the $\boldsymbol{\beta}$ in $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})=0$ for LR.
@@ 3 Technical Details
It is quite time-consuming to compute the Hessian matrix and its inverse in Newton’s method for each iteration. One way to limit this downside is to replace the varying Hessian with a fixed matrix $\bar{H}$. This novel technique is called the fixed Hessian Newton’s method. Böhning and Lindsay (1988) have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\boldsymbol{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})-\bar{H}$ is non-negative definite. With such a fixed Hessian matrix $\bar{H}$, the iteration for Newton’s method can be simplified to:
$\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}-\bar{H}^{-1}\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}).$
Böhning and Lindsay also suggest the fixed matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ is a good lower bound for the Hessian of the log-likelihood function $l(\boldsymbol{\beta})$ in LR.
@@ 3.1 Simplified Fixed Hessian
Bonte and Vercauteren (2018) simplify this lower bound $\bar{H}$ further due to the need for inverting the fixed Hessian in the encrypted domain. They replace the matrix $\bar{H}$ with a diagonal matrix $B$ whose diagonal elements are simply the sums of each row in $\bar{H}$. They also suggest a specific order of calculation to optimize the computation of $B$ more efficiently. Their new approximation $B$ of the fixed Hessian is:
@@
$\boldsymbol{\beta}_{t+1}$ $=\boldsymbol{\beta}_{t}-B^{-1}\cdot\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}),$
$=\boldsymbol{\beta}_{t}-\left[\begin{array}[]{cccc}b_{00}&0&\ldots&0\\ 0&b_{11}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&b_{dd}\\ \end{array}\right]\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]$
$=\boldsymbol{\beta}_{t}-\left[\begin{array}[]{c}b_{00}\cdot\nabla_{0}\\ b_{11}\cdot\nabla_{1}\\ \vdots\\ b_{dd}\cdot\nabla_{d}\\ \end{array}\right],$
where $b_{ii}$ is the reciprocal of $\sum_{i=0}^{d}\bar{h}_{0i}$ and $\nabla_{i}$ is the element of $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})$.
@@
$\boldsymbol{\beta}_{t+1}$ $=\boldsymbol{\beta}_{t}-(-\eta)\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]=\boldsymbol{\beta}_{t}+\eta\cdot\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}),$
@@
The SFH method proposed by Bonte and Vercauteren (2018) has two limitations: (a) in the construction of the simplified fixed Hessian matrix, all entries in the symmetric matrix $\bar{H}$ need to be non-positive. For machine learning applications, datasets are typically normalized in advance to the range [0,1], satisfying the convergence condition of the SFH method. However, in other cases, such as numerical optimization, this condition may not always hold; and (b) the simplified fixed Hessian matrix $B$, as well as the fixed Hessian matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ , can still be singular, especially when the dataset is a high-dimensional sparse matrix, such as the MNIST datasets. We extend their work by removing these limitations so as to generalize this simplified fixed Hessian to be invertible in any case and propose a faster gradient variant, which we term quadratic gradient.

### v8

@@ Abstract
Training logistic regression over encrypted data has been a compelling approach in addressing security concerns for several years. In this paper, we introduce an efficient gradient variant, called $quadratic$ $gradient$, for privacy-preserving logistic regression training. We enhance Nesterov’s Accelerated Gradient (NAG), Adaptive Gradient Algorithm (Adagrad) and Adam algorithms by incorporating their quadratic gradients and evaluate these improved algorithms on various datasets. Experimental results demonstrate that the enhanced algorithms achieve significantly improved convergence speed compared to traditional first-order gradient methods. Moreover, we applied the enhanced NAG method to implement homomorphic logistic regression training, achieving comparable results within just 4 iterations. There is a good chance that the quadratic gradient approach could integrate first-order gradient descent/ascent algorithms with the second-order Newton-Raphson methods, and that it could be applied to a wide range of numerical optimization problems.
@@ 1 Introduction
Given a person’s healthcare data related to a certain disease, we can train a logistic regression (LR) model capable of telling whether or not this person is likely to develop this disease. However, such personal health information is highly private to individuals. The privacy concern, therefore, becomes a major obstacle for individuals to share their biomedical data. The most secure solution is to encrypt the data into ciphertexts first by Homomorphic Encryption (HE) and then securely outsource the ciphertexts to the cloud, without allowing the cloud to access the data directly. iDASH is an annual competition that aims to call for implementing interesting cryptographic schemes in a biological context. Since 2014, iDASH has included the theme of genomics and biomedical privacy. The third track of the 2017 iDASH competition and the second track of the 2018 iDASH competition were both to develop HE-based solutions for building an LR model over encrypted data.
Several studies on logistic regression models are based on homomorphic encryption. Kim et al. Kim et al. (2018b) discussed the problem of performing LR training in an encrypted environment. They used the full batch gradient descent in the training process and the least-squares method to get the approximation of the sigmoid function. In the iDASH 2017 competition, Bonte and Vercauteren Bonte & Vercauteren (2018), Kim et al. Kim et al. (2018a), Chen et al. Chen et al. (2018), and Crawford et al. Crawford et al. (2018) all investigated the same problem that Kim et al. Kim et al. (2018b) studied. In the iDASH competition of 2018, Kim et al. Kim et al. (2019b) and Blatt et al. Blatt et al. (2019) further worked on it for an efficient packing and semi-parallel algorithm. There are other related works (Kim et al., 2019a; Bergamaschi et al., 2019; Ogilvie et al., 2020) focusing on various aspects but the papers most relevant to this work are (Bonte & Vercauteren, 2018) and (Kim et al., 2018a). Bonte and Vercauteren Bonte & Vercauteren (2018) developed a practical algorithm called the simplified fixed Hessian (SFH) method. Our study extends their work and adopts the ciphertext packing technique proposed by Kim et al. Kim et al. (2018a) for efficient homomorphic computation.
@@
    We propose a new gradient variant, quadratic gradient, which can combine the first-order gradient algorithms and the second-order Newton-Raphson method (aka Newton’s method).
@@
    We develop three enhanced gradient algorithms by equipping the original ones with quadratic gradient. The resulting algorithms in most cases demonstrate high performance in terms of convergence speed.
@@
    We implement privacy-preserving logistic regression training using the enhanced NAG method, to our best knowledge, which seems to be a great choice without compromising much on computation and storage.
@@ 2 Preliminaries
We use the square brackets “[][\ ][ ]” to denote the index of a vector or matrix element in what follows. For example, for a vector $\bm{v}\in\mathbb{R}^{(n)}$ and a matrix $M\in\mathbb{R}^{m\times n}$, $\bm{v}[i]$ or $\bm{v}_{[i]}$ means the $i$-th element of vector $\bm{v}$ and $M[i][j]$ or $M_{[i][j]}$ the $j$-th element in the $i$-th row of $M$.
@@ 2.1 Fully Homomorphic Encryption
Fully Homomorphic Encryption (FHE) is a type of cryptographic scheme that can be used to compute an arbitrary number of additions and multiplications directly on the encrypted data. It was not until 2009 that Gentry constructed the first FHE scheme via a bootstrapping operation (Gentry, 2009). FHE schemes themselves are computationally time-consuming; the choice of dataset encoding matters likewise to the efficiency. In addition to these two limits, how to manage the magnitude of plaintext (Jäschke & Armknecht, 2016) also contributes to the slowdown. Cheon et al. Cheon et al. (2017) proposed a method to construct an HE scheme with a rescaling procedure which could eliminate this technical bottleneck effectively. We adopt their open-source implementation HEAAN while implementing our homomorphic LR algorithms. In addition, it is inevitable to pack a vector of multiple plaintexts into a single ciphertext for yielding a better amortized time of homomorphic computation. HEAAN supports a parallel technique (aka SIMD) to pack multiple complex numbers in a single polynomial and provides rotation operation on plaintext slots. The underlying HE scheme in HEAAN is well described in (Kim et al., 2018a; b; Han et al., 2019).
@@ 2.2 Database Encoding Method
Kim et al. Kim et al. (2018a) proposed an efficient and promising database-encoding method by using SIMD technique, which could make full use of the computation and storage resources. Suppose that a database has a training dataset consisting of $n$ samples with $(1+d)$ covariates, they packed the training dataset $Z$ into a single ciphertext in a row-by-row manner.
@@
Han et al. Han et al. (2019) mentioned several basic but important operations used by Kim et al. Kim et al. (2018a) in their implementation, such as a procedure named “SumColVec” to compute the summation of the columns of a matrix. With these fundamental operations, more intricate computations, like computing gradients in logistic regression models, become achievable.
@@ 2.3 Logistic Regression Model
Logistic regression is widely used in binary classification tasks to infer whether a binary-valued variable belongs to a certain class or not. LR can be generalized from linear regression (Murphy, 2012) by mapping the whole real line $(\bm{\beta}^{\top}\mathbf{x})$ to $(0,1)$ via the sigmoid function $\sigma(z)=1/(1+\exp(-z))$, where the vector $\bm{\beta}\in\mathbb{R}^{(1+d)}$ is the main parameter of LR and the vector $\mathbf{x}=(1,x_{1},\ldots,x_{d})\in\mathbb{R}^{(1+d)}$ the input covariate. Thus logistic regression can be formulated with the class label $y\in\{\pm 1\}$ as follows:
$\Pr(y=+1|\mathbf{x},\bm{\beta})$ $=\sigma(\bm{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{-\bm{\beta}^{\top}\mathbf{x}}},$
$\Pr(y=-1|\mathbf{x},\bm{\beta})$ $=1-\sigma(\bm{\beta}^{\top}\mathbf{x})$ $=\frac{1}{1+e^{+\bm{\beta}^{\top}\mathbf{x}}}.$
@@
The logistic regression problem can be transformed into an optimization problem that seeks a parameter $\bm{\beta}$ to maximize $L(\bm{\beta})=\prod_{i=1}^{n}\Pr(y_{i}|\mathbf{x}_{i},\bm{\beta})$ or its log-likelihood function $l(\bm{\beta})$ for convenience in the calculation:
$l(\bm{\beta})=\ln L(\bm{\beta})=-\sum_{i=1}^{n}\ln(1+e^{-y_{i}\bm{\beta}^{\top}\mathbf{x}_{i}}),$
where $n$ is the number of examples in the training dataset. LR does not have a closed form of maximizing $l(\bm{\beta})$ and two main methods are adopted to estimate the parameters of an LR model: (a) gradient descent method via the gradient; and (b) Newton’s method by the Hessian matrix. The gradient and Hessian of the log-likelihood function $l(\bm{\beta})$ are given by, respectively:
$\nabla_{\bm{\beta}}l(\bm{\beta})$ $=\sum_{i}(1-\sigma(y_{i}\bm{\beta}^{\top}\mathbf{x}_{i}))y_{i}\mathbf{x}_{i},$
$\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$ $=\sum_{i}(y_{i}\mathbf{x}_{i})(\sigma(y_{i}\bm{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\bm{\beta}^{\top}\mathbf{x}_{i})(y_{i}\mathbf{x}_{i})$
@@
where $S$ is a diagonal matrix with entries $S_{ii}=(\sigma(y_{i}\bm{\beta}^{\top}\mathbf{x}_{i})-1)\sigma(y_{i}\bm{\beta}^{\top}\mathbf{x}_{i})$ and $X$ the dataset.
The log-likelihood function $l(\bm{\beta})$ of LR has at most a unique global maximum (Allison, 2008), where its gradient is zero. Newton’s method is a second-order technique to numerically find the roots of a real-valued differentiable function, and thus can be used to solve the $\bm{\beta}$ in $\nabla_{\bm{\beta}}l(\bm{\beta})=0$ for LR.
@@ 3 Technical Details
It is quite time-consuming to compute the Hessian matrix and its inverse in Newton’s method for each iteration. One way to limit this downside is to replace the varying Hessian with a fixed matrix $\bar{H}$. This novel technique is called the fixed Hessian Newton’s method. Böhning and Lindsay Böhning & Lindsay (1988) have shown that the convergence of Newton’s method is guaranteed as long as $\bar{H}\leq\nabla_{\bm{\beta}}^{2}l(\bm{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\bm{\beta}$ and $\leq$” denotes the Loewner ordering in the sense that the difference $\nabla_{\bm{\beta}}^{2}l(\bm{\beta})-\bar{H}$ is non-negative definite. With such a fixed Hessian matrix $\bar{H}$, the iteration for Newton’s method can be simplified to:
$\bm{\beta}_{t+1}=\bm{\beta}_{t}-\bar{H}^{-1}\nabla_{\bm{\beta}}l(\bm{\beta}).$
Böhning and Lindsay also suggest the fixed matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ is a good lower bound for the Hessian of the log-likelihood function $l(\bm{\beta})$ in LR.
@@ 3.1 Simplified Fixed Hessian
Bonte and Vercauteren Bonte & Vercauteren (2018) simplify this lower bound $\bar{H}$ further due to the need for inverting the fixed Hessian in the encrypted domain. They replace the matrix $\bar{H}$ with a diagonal matrix $B$ whose diagonal elements are simply the sums of each row in $\bar{H}$. They also suggest a specific order of calculation to optimize the computation of $B$ more efficiently. Their new approximation $B$ of the fixed Hessian is:
@@
$\bm{\beta}_{t+1}$ $=\bm{\beta}_{t}-B^{-1}\cdot\nabla_{\bm{\beta}}l(\bm{\beta}),$
$=\bm{\beta}_{t}-\left[\begin{array}[]{cccc}b_{00}&0&\ldots&0\\ 0&b_{11}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&b_{dd}\\ \end{array}\right]\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]=\bm{\beta}_{t}-\left[\begin{array}[]{c}b_{00}\cdot\nabla_{0}\\ b_{11}\cdot\nabla_{1}\\ \vdots\\ b_{dd}\cdot\nabla_{d}\\ \end{array}\right],$
where $b_{ii}$ is the reciprocal of $\sum_{i=0}^{d}\bar{h}_{0i}$ and $\nabla_{i}$ is the element of $\nabla_{\bm{\beta}}l(\bm{\beta})$.
@@
$\bm{\beta}_{t+1}$ $=\bm{\beta}_{t}-(-\eta)\cdot\left[\begin{array}[]{c}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\\ \end{array}\right]=\bm{\beta}_{t}+\eta\cdot\nabla_{\bm{\beta}}l(\bm{\beta}),$
@@
The SFH method proposed by Bonte and Vercauteren Bonte & Vercauteren (2018) has two limitations: (a) in the construction of the simplified fixed Hessian matrix, all entries in the symmetric matrix $\bar{H}$ need to be non-positive. For machine learning applications, datasets are typically normalized in advance to the range [0,1], satisfying the convergence condition of the SFH method. However, in other cases, such as numerical optimization, this condition may not always hold; and (b) the simplified fixed Hessian matrix $B$, as well as the fixed Hessian matrix $\bar{H}=-\frac{1}{4}X^{\top}X$ , can still be singular, especially when the dataset is a high-dimensional sparse matrix, such as the MNIST datasets. We extend SFH by removing these limitations to generalize this simplified fixed Hessian to be invertible in any case and propose a faster gradient variant, which we term $quadratic$ $gradient$.